6 September 2013

Configured for LBC

02:00 Starting with Flat Fields. Skies photometric.

Uspec (021016...) and V-BESSEL (021013)
  • in 1 sec 8000 in b and 6000 in r...
  • try 3 sec. Still low counts... 11k and 8k. * b 021443 Uspec has a dark splotch near the top right of chip 1. Will take
  • Will try other set of filters, but afraid it is already too dark.
    • both V and r low (1600 and 1300, respectively), but texp=1 sec only. Upped time to 10sec.
    • now r and V on B and Red with 30-sec.
      • These tests are now not to collect flats - unfortunately I started too late - but to check out illumination through filters.

  • On ACT0302. Running dofpia (but I forgot the first... will see how it goes anyway).
    • Finds lots of pupils. Red initially have negative Z11, Z22, but blue look pretty uniformly illuminated, albeit too large so only Z4 is sent.
    • Red came close to convergence in two iterations and est seeing 0.8 (just like the DIMM... wow).
    • Blue - on 3rd iteration is close to convergence, with est seeing 1.1.
    • Red/Blue converged on 4 iterations. Est seeing ~1" red and 1.8" blue. Yes, the blue pupils still have some extension like the skirt.
  • Run pointing star OB for ACT0302
    • 024405 b and 024404 r
Sending the calculated optics offsets:
SX: dX = 0.063 dY = -0.038 dRX = -0.824 dRY = -1.342
DX: dX = -0.071 dY = 0.007 dRX = 0.147 dRY = 1.516
    • repeat

Tilt Data
  • On NGC6779_offset. Airmass = 1.01. El=81 deg.
  • PA=0, RB_rV_superfoc7. Blue images listed below...
 
lbcb.20130906.025648.fits[2] RB_rV_superfoc7
lbcb.20130906.025737.fits[2] RB_rV_superfoc7
lbcb.20130906.025827.fits[2] RB_rV_superfoc7
lbcb.20130906.025916.fits[2] RB_rV_superfoc7
lbcb.20130906.030005.fits[2] RB_rV_superfoc7
lbcb.20130906.030055.fits[2] RB_rV_superfoc7
lbcb.20130906.030144.fits[2] RB_rV_superfoc7
  • PA=180, RB_rV_superfoc7_pa180
    • LBC crashed during this. I only saw a popup saying that it would automatically restart. 03:08 UT approx. Looking for error... looks like it was a blue camera error,but was preceded by a possibly unrelated blue rotator warning.
2013/09/06 03:06:21.455698 W B                   rotator system recovered from a problem [src/channel/channel.c:1132]
2013/09/06 03:06:21.455751 N B                   preset completed
2013/09/06 03:06:21.455781 E B                   >>> preset/expose/save interrupted at preset time

then later...

2013/09/06 03:06:21.455781 E B                   >>> preset/expose/save interrupted at preset time
2013/09/06 03:08:23.152381 E B CAMERA            RPC Client/Host communication failed retval:3 [src/camera/camera.c:1101]

Got 4 blue images before crash.
lbcb.20130906.030359.fits[2] RB_rV_superfoc7_pa
lbcb.20130906.030449.fits[2] RB_rV_superfoc7_pa
lbcb.20130906.030538.fits[2] RB_rV_superfoc7_pa
lbcb.20130906.030642.fits[2] RB_rV_superfoc7_pa

03:20 dofpia after crash - on same NGC6779 field. Now at 86 deg elevation so will run superfoc 2x here, but move to a field at lower elevation.
  • dofpia converged with est seeing 0.83" on R, est 1.4 on B. DIMM last measurement is 0.8" at 3:14 UT.
lbcb.20130906.032232.fits[2] RB_rV_superfoc7_pa
lbcb.20130906.032321.fits[2] RB_rV_superfoc7_pa
lbcb.20130906.032409.fits[2] RB_rV_superfoc7_pa
lbcb.20130906.032458.fits[2] RB_rV_superfoc7_pa
lbcb.20130906.032547.fits[2] RB_rV_superfoc7_pa
lbcb.20130906.032637.fits[2] RB_rV_superfoc7_pa
During PA=0 sequence LBC crashed.
2013/09/06 03:33:23.190290 E R CAMERA            RPC Client/Host communication failed retval:3 [src/camera/camera.c:1101]
LBC restarted automatically. As this field is at 87 deg elev now, we'll go to NGC110 for now, then probably back to catch NGC6779 as it it is setting.

Note by KS: both of these crashes were core-dumps by the lbc software. It should have recovered from the RPC failures more gracefully than shutting everything down and coming back up. The core-dump fix was put in at 03:57UTC.

03:42 David going to star near NGC110, pointing star ACT0400. Airmass 1.8. Elev=33 deg. Note that corrections to Z4, Z11 and Z22 are -ve, +ve, +ve. Converged in 3 iterations, with est seeing = 1.39 R and 1.88 B.

03:50 Start PA=0 superfoc7 here.
  • arrghhh ... on the ACT400 star...
  • run lbcrangebal:
MOUNT: delta_IE = 9.22 arcsec, delta_CA = -7.66 arcsec
M1 SX: dX = 0.11 mm, dY = 0.08 mm, dRX = 1.80 arcsec, dRY = -2.36 arcsec 
M1 DX: dX = -0.17 mm, dY = -0.10 mm, dRX = -2.14 arcsec, dRY = 3.69 arcsec 

03:56 On NGC110. dofpia.
  • Converged in 1 iteration with est seeing 1.12 R and 1.52 B. Seeing not really good enough for tilt measurement, but will see if we have RPC errors.
  • PA=0.
bcb.20130906.035928.fits[2] RB_rV_superfoc7
lbcb.20130906.040016.fits[2] RB_rV_superfoc7
lbcb.20130906.040106.fits[2] RB_rV_superfoc7
lbcb.20130906.040156.fits[2] RB_rV_superfoc7
lbcb.20130906.040245.fits[2] RB_rV_superfoc7
lbcb.20130906.040334.fits[2] RB_rV_superfoc7
lbcb.20130906.040422.fits[2] RB_rV_superfoc7
  • PA=180
lbcb.20130906.040842.fits[2] RB_rV_superfoc7_pa
lbcb.20130906.040930.fits[2] RB_rV_superfoc7_pa
lbcb.20130906.041019.fits[2] RB_rV_superfoc7_pa
lbcb.20130906.041107.fits[2] RB_rV_superfoc7_pa
lbcb.20130906.041156.fits[2] RB_rV_superfoc7_pa
lbcb.20130906.041245.fits[2] RB_rV_superfoc7_pa
lbcb.20130906.041335.fits[2] RB_rV_superfoc7_pa

04:16 Odd behavior... I started to run 'dofpia' but then, deciding to go back to the higher elevation field, stopped it. But I mistakenly clicked the "stop" button on the GUI - then I did a cntl-c in the dofpia window. All of the buttons - upload, play, pause, stop - were 'greyed out'. And HK and all systems were enabled. I did not see any popup box this time. These are the two lines in the log file before the log service was restarted. It looks like an RPC error:
2013/09/06 04:14:52.439565 N R ROTATOR           moving from +80.877 deg to -99.583 deg position
2013/09/06 04:14:58.389758 E R CAMERA            RPC Client/Host communication failed retval:3 [src/camera/camera.c:1101]
    Note by KS: this is the same core-dump as the previous two. It was fixed at 03:57UTC, so the restart after this used the new version.

04:26 Slewing back to ACT0302 for NGC6779 field.
  • dofpia, /X2
    • Blue converged in 2 iterations, est seeing 1.27 and Red converged in 3 iterations with est seeing 0.95".
  • lbcrangebal: B 043457, R 043453
  • Played again but stopped the Pointing OB - I get an error box: "OB execution error: right main mirror failed to focus! Execution is continuing!" Is the older version of the software in place? I thought this had been fixed?
    Note by KS: This is a different piece of the code, not the standard channel, filter, rotator, camera, tracker that I changed to check for STOPPED. Found the focus code and hopefully this is fixed now.
MOUNT: delta_IE = 13.88 arcsec, delta_CA = 4.53 arcsec
M1 SX: dX = -0.12 mm, dY = 0.02 mm, dRX = 0.43 arcsec, dRY = 2.52 arcsec 
M1 DX: dX = 0.25 mm, dY = 0.20 mm, dRX = 4.19 arcsec, dRY = -5.37 arcsec 

04:38 Slewed to NGC6779_offset. When I played and stopped this OB, no error box appeared.
  • dofpia, /X2
  • Converged in 1 iter with seeing on Blue 1.6" and Red 1.4". Pupils look less well defined.
  • repeat dofpia, /X2 while David gets a DIMM seeing measurement. 1.4" on the DIMM
    • converged in 2 iter est seeing 1.09 R and 1.4 Blue.
  • PA=0 superfoc (El = 74, airmass = 1.04)
  • PA=180 superfoc (El = 71). DIMM 2.1"

05:02 DIMM seeing 2.1 at El 71 and David says there are clouds coming. We'll do the test which John requested - running the OB with large dithers to see if there is rotational trailing. First with the old config file and then with the config file changed.
  • Copied OB for NGC6779_offset to Calib_OBs/TCStesting as NGC6779_offset_rottrail and edited to use just one filter per channel (using V-BESSEL for both Blue and Red) and exptime=180sec. Edited dither pattern to be what it was when the problem was seen:
dithoffx = -700, -350, 0, 350, 700 and dithoffy = -140, -70, 0, 70, 140 
  • Use current configuration files, in /home/lbccontrol/conf/ blue/redchannel.conf which have the value of GETROTATOR TRAJECTORY FROM TELESCOPE = TRUE: Sequence starts with 051531 b and 051521 r
    • (before I had failed to change the number of filters and exptime --- wanted to increase exptime to see trailing --- seeing estimate from these images is 1.7") DIMM during 0515xx pair is 2". DIMM star coming and going due to clouds.
    • in fact, lost guiding during some exposures.
      • During the exposure, guiding reports failed and working. I glanced at it once and saw a "?" for both red and blue thumbnail and the Guiding Sts: 8.0s working. But soon after it reported "failed", so I think the thumbnail and Sts are a little out of sync... just takes my getting used to watching it.

05:30 Now changed both bluechannel.conf and redchannel.conf files (in /home/lbccontrol/conf) to set GETROTATOR TRAJECTORY FROM TELESCOPE = FALSE
cd into /home/lbccontrol/conf:
[root@CMU conf]# cp bluechannel.conf bluechannel_orig_20130906.conf
[root@CMU conf]# cp redchannel.conf redchannel_orig_20130906.conf
[root@CMU conf]# vi redchannel.conf
[root@CMU conf]# vi bluechannel.conf
[root@CMU conf]# grep Channel.GetRotatorTrajectoryFromTelescope *channel.conf
bluechannel.conf:Channel.GetRotatorTrajectoryFromTelescope = FALSE           # TRUE or FALSE to enable telescope downloading of rotator trajectory or LBC calculation
redchannel.conf:Channel.GetRotatorTrajectoryFromTelescope = FALSE           # TRUE or FALSE to enable telescope downloading of rotator trajectory or LBC calculation
Now run lbckill/lbcstart

05:40 Run dofpia,/X2 on field before repeating rottrail ob.
  • DIMM seeing is 1.4" In fact dofpia est seeing does look better too, 1.1" in R and 1.4" in B.. El = 62 now.
  • Play NGC6779_offset_rottrail OB again now using internal rotator trajectory.
    • first images are 054829 b and 054825 r. Again guiding is off and on.
  • Note --- On the gui I see Z4 always 0.0. However I see in the lbc.log that all P03 values are 0.0. There are pupil images on the active optics tech chip (_2) --- is there a problem with tech chip fpia? Maybe there is not a good isolated pupil, or because some of the images show no pupils due to clouds (would the program then reject the whole set?), the computation cannot be made?
  • Seeing is blowing up again - 2.1" on blue and 1.8" on red.

06:10 Restore lbc conf files:
[root@CMU conf]# cp redchannel_orig_20130906.conf redchannel.conf
cp: overwrite `redchannel.conf'? y
[root@CMU conf]# cp bluechannel_orig_20130906.conf bluechannel.conf
cp: overwrite `bluechannel.conf'? y
[root@CMU conf]# grep Channel.GetRotatorTrajectoryFromTelescope *channel.conf
bluechannel.conf:Channel.GetRotatorTrajectoryFromTelescope = TRUE            # TRUE or FALSE to enable telescope downloading of rotator trajectory or LBC calculation
redchannel.conf:Channel.GetRotatorTrajectoryFromTelescope = TRUE            # TRUE or FALSE to enable telescope downloading of rotator trajectory or LBC calculation
lbckill/lbcstart

06:15 Closed for clouds.

While closed, Dave ran some LUCI calibrations and I ran some MODS calibrations. Dome dark, telescope at zenith.

MODS: grlamps, grflats, prlamps. prflats (hung on 5th blue exposure?) Will repeat prflats another time.

08:40 Re-open. LBCs still.

08:44 Slew to ACT400. dofpia,/first,/X2. Both blue and red pupil images appear to be nearly in focus?? dofpia was doing the right thing on blue at least, on red the last 'pupil' image looked extremely comatic(!? -- see lbcr.20130906.085020) but stopped it to ask David to clear active optics.
  • AHA - we were no longer connected to the telescope. dofpia was not "playing" the OB - the "play" button was black. Reconnected LBC to telescope and repeated. Not sure how LBC became disconnected.
    Note by KS: The TCSON command started at 06:12:29 after the restart when you restored the config files. But, we never see the completion of the command in the log file. Don't know what happened there, but you're right it should have been connected when you went back to LBC at 08:40.
  • first pupils are b 85649 and r 85645. Converged in 3 iterations with est seeing 1.4" on both sides --- not much larger on blue than on red as we normally see.

09:03 ACT400 for lbcrangebal.

09:07 Slew to NGC110. Stop OB - this time I got a popup error box.
  • dofpia,/X2
  • converged in 1 iter, seeing ~ 1.3"
  • PA=0 superfoc7 rV 091201/091206 for R/B
  • PA=0 superfoc7 ig
  • PA=0 superfoc7 rV 092512/092535 for R/B
  • Seeing on DIMM is 1.7" now --- contemplating just switching to LUCI to get something. Repeating the rV/gi/rV is necessary but not in this seeing.

09:33 Reconfiguring for LUCI.
  • luci.20130906.NNNN.fits

10:03 BS9188 w/off-aixs guide star
  • dice5_10x (will stop after ~5 cycles)
  • 5 x 4.0s DCR clear+Ks N375 FC-on
  • Files: 0001 - 0025

10:32 BS9105
  • Files 0026 0027 image pair N375
  • Files 0028 0029 image pair N180
  • ROTCEN N180 (0030=test; FC not ON)
    • 0031 PA=0
    • 0032 180
    • 0033 270
    • 0034 90

10:52 BS9107
  • Files 0035 0036 image pair N180 (denser field than BS9105)
  • ROTCEN N180 (0037=test, FC not ON)
    • 0038 PA=0
    • 0039 180
    • 0040 270
    • 0041 90 (did I forget to take this one?)
  • slitdither.acq 0042=src 0043=slit 0044=src+slit
  • slitdither.sci 0045 - 0054 (aborted script, not guiding at top position)
  • send ABS 0,0 offset 0055, 0056, send REL 0.42,0.99 to re-center 0057
  • slitdither_30s.sci 0058 - 0077 (both positions guided)
  • qualitatively looks good, maybe 1 pixel misalignment in 60" dither along slit.

11:56 End of night. Blind mask and filters, N375, mirror

Barring surprises in the quantitative results, LUCI appears to be ready for science

12:03 Close dome

-- OlgaKuhn - 06 Sep 2013
Topic revision: r6 - 09 May 2014, KelleeSummers
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback